Three people check into a hotel. They find out that if one more person checked in with them, the check in price will be reduced by $50. How much is the Check in price?
Answer by mathslover(132)

let the check in price be x
if 3 people check in then price per person =x/3
if 4 people check in then price per person=x/4
now x/3 - x/4 = 50
x/12 = 50
x=600